three man are seeking public office. candidates a and b are given about the same chance of winning, but candidate c is given twice the chance of either a or b.
a) what is the probability that c wins?
b) what is the probability that a does not win?

Answer by stanbon(75887)

three man are seeking public office. candidates a and b are given about the same chance of winning, but candidate c is given twice the chance of either a or b.
----
P(a) = x
P(b) = x
P(c) = 2x
---
x + x + 2x = 1
4x = 1
x = 1/4
2x = 1/2
----------
a) what is the probability that c wins? = 1/2
----------
b) what is the probability that a does not win?
1-1/4 = 3/4
===============================================
Cheers,
Stan H.
===============